OpenStudy (bluemoon1):

Need help with a couple of questions. Give a couple of minutes to type(:

OpenStudy (bluemoon1):

Find the slope of a line parallel to each given line. 1. x-4y = 4 2. 2x + 5y = 0

OpenStudy (bluemoon1):

@AloneS

alones (alones):

Lol im here

OpenStudy (bluemoon1):

If you know how to do these please also show work(; and I'll medal you plenty

OpenStudy (zyi6):

Parallel lines have the same slope, with that being said, we can put each equation into slope intercept form to see the slope: We will start with the first equation: x - 4y = 4 In order to put this into slope intercept form, we need to subtract x from both sides: -4y = -x + 4 Now, we need to isolate the y by dividing by -4: y = x - 1 So, this is our first equation. Can you do the same thing for the next equation?